Sally bought three chocolate bars and a pack of gum and paid $1.75. Jake bought two chocolate bars and four packs of gum and paid $2.00. Find the cost of a chocolate bar and the cost of a pack of gum.
Answer by checkley75(3666)

3C+G=175 OOR G=175-3C
2C+4G=200 NOW SUBSTITUTE (175-3C) FOR G IN THE SECOND EQUATION & SOLVE FOR C
2C+4(175-3C)=200
2C+700-12C=200
-10C=-500
C=-500/-10
C=50 COST OF THE CHOCOLATE BAR
G=175-3*50
G=175-150
G=25 COST OF THE GUM
PROOF
2*50+4*25=200
100+100=200
200=200